Rewrite the equation 2x-8y=48 as a function of x.
Answer by tutorcecilia(2152)

2x-8y=48 [Solve for y]
-8y=-2x+48
-8y/-8=-2x/-8+(48/-8)
y=1/4x-6
f(x)=1/4x-6